Suppose the labor supply curve for school teachers is given by Ls=100+3W and the labor demand...

  1. Suppose the labor supply curve for school teachers is given by Ls=100+3W and the labor demand curve for school teachers is given by LD=700-W. What is the equilibrium wage and quantity of school teachers?
Answer #1

Answer : Labor supply :

L = 100 + 3W

=> L - 100 = 3W

=> W = (L - 100) / 3

Labor demand :

L = 700 - W

=> W = 700 - L

At equilibrium Demand = Supply occur. So,

700 - L = (L - 100) / 3

=> (700 - L) * 3 = L - 100

=> 2100 - 3L = L - 100

=> 2100 + 100 = L + 3L

=> 4L = 2200

=> L = 2200 / 4

=> L = 550

From demand function we get,

W = 700 - 550

=> W = 150

Therefore, here the equilibrium wage rate is $150 and the quantity of school teacher is 550.